# Exploring the Adversarial Vulnerabilities of Vision-Language-Action Models in Robotics

### v2

@@ Abstract
Recently in robotics, Vision-Language-Action (VLA) models have emerged as a transformative approach, enabling robots to execute complex tasks by integrating visual and linguistic inputs within an end-to-end learning framework. While VLA models offer significant capabilities, they also introduce new attack surfaces, making them vulnerable to adversarial attacks. With these vulnerabilities largely unexplored, this paper systematically quantifies the robustness of VLA-based robotic systems. Recognizing the unique demands of robotic execution, our attack objectives target the inherent spatial and functional characteristics of robotic systems. In particular, we introduce an untargeted position-aware attack objective that leverages spatial foundations to destabilize robotic actions, and a targeted attack objective that manipulates the robotic trajectory. Additionally, we design an adversarial patch generation approach that places a small, colorful patch within the camera’s view, effectively executing the attack in both digital and physical environments. Our evaluation reveals a marked degradation in task success rates, with up to a 100% reduction across a suite of simulated robotic tasks, highlighting critical security gaps in current VLA architectures. By unveiling these vulnerabilities and proposing actionable evaluation metrics, this work advances both the understanding and enhancement of safety for VLA-based robotic systems, underscoring the necessity for developing robust defense strategies prior to physical-world deployments.
@@ 1 Introduction
Figure 1: Fatal failure rates resulting from different attacking schemes (i.e., UADA, UPA, and TMA) underscore vulnerabilities in VLA-based robotic systems.
In the movie Finch, set in a post-apocalyptic future, an intelligent robot navigates complex themes, underscoring the importance of interactive safety with its human master. Once speculative, this portrayal is now increasingly plausible with the rise of visually-conditioned language models (VLMs) [73, 82] capable of seamlessly interpreting both visual and linguistic contexts. A notable realization of this potential can be seen in Vision-Language Action (VLA) models [76, 16, 35, 77], which integrate VLMs into robotic systems to enable end-to-end learning that encompasses both high-level planning and the generation of low-level robot control actions. While VLA models demonstrate a promising step toward generalist robotic intelligence, they also introduce potential vulnerabilities that remain underexplored. This work addresses the pressing need to investigate the new attack surfaces associated with VLA-based robotic systems.
@@
To address these challenges, our work intensifies the adversarial threats posed to VLA-based systems by developing specialized attack objectives and designing effective attack methods. Specifically, we formulate an Action Discrepancy Objective aimed at maximizing the normalized action discrepancy within VLA-based robotic systems. This strategy ensures that, at each decision point, the robot’s behavior can diverge from the optimal trajectory. Additionally, we introduce Geometry-Aware Objective that considers the robot’s movement in three-dimensional space, characterized by three degrees of freedom. By optimizing the cosine similarity between adversarial and ground-truth directions, we induce deviations in the robot’s movement direction from its intended path, increasing the likelihood of task failure. To achieve these attack objectives, we develop a straightforward yet effective Patch-Based Attack targeting VLA-based robotic systems. This approach enables adversarial attacks in both digital and physical settings, revealing substantial vulnerabilities within the VLA-based system.
Together, these innovations yield several significant contributions: This work presents a pioneering and comprehensive analysis of vulnerabilities within VLA-based robotic systems, a new paradigm for training generalized robotic policies using generative foundation models. We reveal critical threats to adversarial attacks (see Fig. 1), emphasizing the urgent need to enhance robustness before real-world deployment. We are the first to define attack objectives specific to the powerful VLA model and to employ a straightforward adversarial patch against it (see §3). This offers valuable insights for the research community to explore systemic failures in similar concurrent generative foundation models. We rigorously evaluate our approach in both simulated and physical environments across four distinct robotic tasks, observing a significant raises in task failure rates of 100% and 43%, respectively. This highlights the effectiveness of our attack strategies (see §4.2).
@@ 2.1 AI-driven Generalist Robot
Developing generalist robots [63, 6, 38, 27, 14] requires models not only handle varied interactions but also maintain robustness in unstructured environments. Early intelligent robots [57, 53, 54, 56, 86] relied on rule-based approaches, effective only in controlled settings and requiring extensive reprogramming for new tasks. The advent of deep learning [42, 59, 19, 31, 33] and reinforcement learning [66, 75, 1] has shifted this paradigm to adapt through data, increasing robot versatility and reducing manual reconfiguration.
Recent studies [7, 55, 60] have further advanced the development and potential of large-scale visually-conditioned language models (VLMs) as key enablers for generalist robots, demonstrating promising generalization across a variety of scenes [87, 32, 67, 25, 43, 15]. One of the notable examples is OpenVLA [35], which integrates dual visual encoders with a pre-trained large language model to enable observation-to-action capabilities through instruction-following [47, 48, 34, 10, 85, 73]. By aligning visual and textual semantics, OpenVLA facilitates complex scenario understanding and end-to-end action generation, showing impressive generalization when executing previously unseen task instructions. Despite its success in generalist robotic applications, the vulnerability of such systems remains largely unexplored. This study aims to address this gap by investigating adversarial threats to VLA-based robotic systems.
2.2 Human-Robot Interaction Safety
2.3 Adversarial Attacks in Robotics
@@ 3 Methodology
This section outlines our proposed methodology. We first review the core algorithmic principles of VLA in §3.1, which serves as the foundation for the adversarial scheme designs. We then present two types of untargeted adversarial attacks, including Untargeted Action Discrepancy Attack (UADA) in §3.2 and Untargeted Position-aware Attack (UPA) in §3.3. In §3.4, we further include Targeted Manipulation Attack (TMA), designed to direct the robot toward specific erroneous execution. The implementation details are presented in §3.5. The overall framework is shown in Fig. 2.
@@ 3.1 Preliminary
Vision-Language Action models [87, 35] (see Fig. 2) are built on large language models integrated with visual encoders, enabling robots to interpret human instructions and process visual input from a camera to perform context-aware actions. VLA models [35, 87, 44] generally abstract continuous action prediction into a classification problem by discretizing robot actions. Specifically, they first discretize the continuous probabilities into class labels $c=\arg\max F(x)$. An action de-tokenizer then generates actions $y=DT(c)$. By categorizing action values into discrete class labels, the model converts continuous probability outputs into discrete signals, this simplification facilitates quicker convergence and faster training times, and is commonly used for VLA-based models [35, 87, 44].
In this work, we focus on a 7-degree-of-freedom (DoF) robotic arm [23]. At each step, an action consists of 7DoFs with specific physical significance in three-dimensional Cartesian space, represented by:
@@
where $\Delta{P_{x,y,z}}$ and $\Delta{R_{x,y,z}}$ denote relative positional and rotational changes along the x-, y-, and z-axis, discretized into 256 uniform bins across the action range [35]. The $gripper$ state is binary, indicating its open or closed state. This control design presents a unique challenge for adversarial attacks, as finely divided bins result in minimal action discrepancies between neighboring bins (e.g., $\pm$0.007/bin). This means that even if an attack shifts the classification to an adjacent bin, the resulting action discrepancy has minimal impact real-world performance.
@@ 3.2 Untargeted Action Discrepancy Attack
We first focus on maximizing action discrepancies by introducing Untargeted Action Discrepancy Attack (UADA). This attack is based on the observation that larger robot actions usually correlate with intense physical movements, which, in turn, may amplify the potential for real-world hazards [28, 30, 29]. Specifically to UADA, the attack target is one or combination of 7DoFs, defind as $y^{I}=\{y^{i}|i\in[1,\dots,7]\}$. To define UADA’s objective, we first determine the maximum action discrepancy $d_{\text{max}}$, which represents the largest deviation that can occur within the allowed range for each DoF. This is calculated by measuring the distance between the ground truth $y$ and its action range $[y_{\min},y_{\max}]$ as:
$\small d_{max}^{i}(y)=\max(|y^{i}-y^{i}_{\min}|,|y^{i}-y^{i}_{\max}|).$ (2)
We next calculate the applied action discrepancy $d_{\text{applied}}^{i}(x,y)$ $=|F(x)^{i}-y^{i}|$ in order to measure the deviation between the model’s output and the ground truth, where $F(x)^{i}$ is the $i$-th action output of model $F(x)$. In this sense, we can evaluate the Normalized Action Discrepancy (NAD) as the drifting ratio between $d_{applied}^{i}$ and $d_{max}^{i}$ as:
$\rm{NAD}=\frac{1}{I}\sum_{i}^{I}d^{i}_{\text{applied}}/d^{i}_{\text{max}}.$ (3)
Since VLA models are currently operated via predicting discrete token labels (see §3.1), we then need to bridge action discrepancy derived from these discrete labels into learnable targets. Specifically, we weighted sum the model output probabilities across $J$ bins by their own class label $j$ as:
$d_{\text{applied}}^{i\prime}(x,y)=|\sum_{j=1}^{J}(j\cdot\mathcal{F}(x)_{j}^{i})-c^{i}|,$ (4)
where $\mathcal{F}(\cdot)_{j}^{i}$ is the model output probability of $i$-th DoF’s $j$-th bin. Finally, the attack objective for UADA is defined as minimizing the ratio of weighted sum $d^{\prime}_{\text{applied}}$ to $d_{max}(y^{i})$:
$\small\mathcal{L}_{\text{UADA}}=\mathbb{E}_{(x,y)\sim\mathcal{X}}\left[\sum^{I}_{i}\frac{d^{i\prime}_{\text{applied}}(x+\delta,y)}{d^{i}_{\max}(y)}\right]^{-1},$ (5)
where $x,y$ is a benign image and action pair, sampled from distribution $\mathcal{X}$. UADA defines the objective functions tailored for the robot’s action space, allowing adversarial patches to create significant, lasting disruptions in task performance. In this design, we uniquely leverage the model’s discrete action structure to design attacks that amplify meaningful discrepancies in robot behavior.
@@ 3.3 Untargeted Position-aware Attack
We further extend untargeted adversarial attacks to consider the positional dynamics within VLA models. In typical task execution, precise and directed movements of the end-effector towards designated goals are essential, with actions mapped in three-dimensional space. The positional degrees of freedom, $y^{pos}=[y^{1},y^{2},y^{3}]$ encapsulates the directional movements within 3D for effective robotic control.
Recognizing the importance of $y^{pos}$ in controlling the end-effector’s path, we introduce a position-aware attack to disrupt the intended movement trajectory. This type of adversarial vulnerability remains largely unexplored but has significant implications for task failure: the attack objective can steer the end-effector away from its intended path by introducing directional perturbations, amplifying errors and causing substantial trajectory distortions. ${}_{\!}$we define the Untargeted Position-aware Attack (UPA) objective as:
$\small\mathcal{L}_{\text{UPA}}=\mathbb{E}_{(x,y)\sim\mathcal{X}}\big{[}\alpha\frac{y^{pos}_{adv}\cdot y^{pos}}{\|y^{pos}_{adv}\|\|y^{pos}\|}+\beta\frac{1}{||y^{pos}_{adv}-y^{pos}||_{2}}\big{]},$ (6)
@@
1:  Input: $\mathcal{X}$: dataset; $\delta$: adversarial patch; $\mathcal{L}_{o}$: objective function; $\mathcal{F}$: vision language action model; $\mathcal{T}(\cdot)$: Transfromation; $\phi,\psi$ are two hyperparameters for transformation; T: maximum iteration steps; $k$: maximum inner-loop steps.

### v3

@@ Abstract
Recently in robotics, Vision-Language-Action (VLA) models have emerged as a transformative approach, enabling robots to execute complex tasks by integrating visual and linguistic inputs within an end-to-end learning framework. While VLA models offer significant capabilities, they also introduce new attack surfaces, making them vulnerable to adversarial attacks. With these vulnerabilities largely unexplored, this paper systematically quantifies the robustness of VLA-based robotic systems. Recognizing the unique demands of robotic execution, our attack objectives target the inherent spatial and functional characteristics of robotic systems. In particular, we introduce two untargeted attack objectives that leverage spatial foundations to destabilize robotic actions, and a targeted attack objective that manipulates the robotic trajectory. Additionally, we design an adversarial patch generation approach that places a small, colorful patch within the camera’s view, effectively executing the attack in both digital and physical environments. Our evaluation reveals a marked degradation in task success rates, with up to a 100% reduction across a suite of simulated robotic tasks, highlighting critical security gaps in current VLA architectures. By unveiling these vulnerabilities and proposing actionable evaluation metrics, we advance both the understanding and enhancement of safety for VLA-based robotic systems, underscoring the necessity for continuously developing robust defense strategies prior to physical-world deployments [^1] Our code is available at https://github.com/William-wAng618/roboticAttack. Project Homepage at https://vlaattacker.github.io/.
@@ 1 Introduction
Figure 1: Adversarial Vulnerabilities induced by malicious manipulation. (A). Illustration of adversarial threats in robotic task execution. (B). Example of semantic-rich adversarial patches generated by proposed methods. (C). Comparison of failure rates across different attack schemes (UADA, UPA, and TMA).
In the movie Finch, set in a post-apocalyptic future, an intelligent robot navigates complex themes, underscoring the importance of interactive safety with its human master. Once speculative, this portrayal is now increasingly plausible with the rise of Large Vision Language Models (LVLMs) [70, 79], capable of seamlessly interpreting both visual and linguistic contexts. A notable realization of this potential can be seen in Vision-Language Action (VLA) models [73, 16, 34, 74], which integrate LVLMs into robotic systems to enable end-to-end learning that encompasses both high-level trajectory planning and low-level robot action control. While VLA models demonstrate a promising step toward generalist robotic intelligence, they also introduce potential vulnerabilities that remain largely underexplored. This work, in light of this view, pioneers the understandings of the pressing need to investigate the new attack surfaces associated with VLA-based robotic systems.
@@
To address these challenges, our work intensifies the adversarial threats posed to VLA-based systems by both developing specialized attack objectives and designing effective attack methods. Specifically, we formulate an Action Discrepancy Objective aimed at maximizing the action discrepancy within VLA-based robotic systems. This strategy ensures that, at each decision point, the robot’s behavior can diverge from the optimal trajectory. Additionally, we introduce Geometry-Aware Objective that considers the robot’s movement in three-dimensional space, characterized by three degrees of freedom. By optimizing the cosine similarity between adversarial and ground-truth directions, we induce deviations in the robot’s movement direction from its intended path, increasing the likelihood of task failure. To achieve these attack objectives, we develop straightforward yet effective Patch-Based Attacks targeting VLA-based robotic systems. This approach enables adversarial attacks in both digital and physical settings, revealing substantial vulnerabilities within the VLA-based system.
Altogether, these innovations yield several significant contributions: This work presents a pioneering and comprehensive analysis of vulnerabilities within VLA-based robotic systems, a new paradigm for training generalized robotic policies using generative foundation models. We reveal critical threats to adversarial attacks, emphasizing the urgent need to enhance robustness before real-world deployment; To the best of our knowledge, we are the first work to define attack objectives specific to the powerful VLA models and to employ a straightforward adversarial patch against it (see §3). This offers valuable insights for the research community to explore systemic failures in similar concurrent generative foundation models; We rigorously evaluate our approach in both simulated and real-world environments across four distinct robotic tasks, observing a significant raises in task failure rates of 100% and 43%, respectively. This highlights the effectiveness of our attack strategies (see §4.3).
@@ 2.1 AI-driven Generalist Robot
Developing generalist robots [60, 6, 37, 27, 14] requires models to not only handle varied interactions but also maintain robustness in unstructured environments. Early intelligent robots [55, 51, 52, 54, 83] generally relied on rule-based approaches, effective only in controlled settings and needing extensive reprogramming for new tasks. The advent of deep learning [40, 56, 19, 31] and reinforcement learning [63, 72, 1] has shifted this paradigm to adapt through the data-driven pipeline, increasing robot versatility and reducing manual reconfiguration.
Recent studies [7, 53, 57] have further advanced the development and potential of Large Vision Language Models (LVLMs) as key enablers for generalist robots, demonstrating promising generalization across a variety of scenes [84, 32, 64, 25, 41, 15]. One of the notable examples is OpenVLA [34], which integrates dual visual encoders with a pre-trained large language model to enable observation-to-action capabilities through instruction-following [45, 46, 33, 10, 82, 70]. By aligning visual and textual semantics, OpenVLA facilitates complex scenario understanding and end-to-end action generation, showing impressive generalization when executing previously unseen task instructions.
2.2 Adversarial Attacks in Robot
@@ 3 Methodology
In this section, we first review the core algorithmic principles of VLA in §3.1, which serves as the foundation for the adversarial scheme designs. We then present two types of untargeted adversarial attacks, including Untargeted Action Discrepancy Attack (UADA) in §3.2 and Untargeted Position-aware Attack (UPA) in §3.3. In §3.4, we further include Targeted Manipulation Attack (TMA), designed to direct the robot toward specific erroneous execution. Finally, we introduce the Normalized Action Discrepancy (NAD) metric in §3.5. The overall framework is shown in Fig. 2.
@@ 3.1 Preliminary
Vision-Language Action models [84, 34] (see Fig. 2) are built on large language models integrated with visual encoders, enabling robots to interpret human instructions and process visual input from a camera to perform context-aware actions. VLA models [34, 84, 42] generally abstract continuous action predictions into a classification problem by discretizing robot actions. Specifically, they first discretize the continuous probabilities into class labels $y=\arg\max\mathcal{F}(x)$ where $\mathcal{F(\cdot)}$ is the VLA model. An action de-tokenizer then generates actions $A=DT(y)$. By categorizing action values into discrete class labels, the model converts continuous probability outputs into discrete signals, this simplification facilitates quicker convergence and faster training times, and is commonly used for VLA-based models [34, 84, 42].
In this work, we focus on a 7 degree-of-freedoms (DoFs) robotic arm [23]. At each step, an action consists of 7 DoFs with specific physical significance in three-dimensional Cartesian space, represented by:
@@
where $\Delta{P_{x,y,z}}$ and $\Delta{R_{x,y,z}}$ denote relative positional and rotational changes along the x-, y-, and z-axis, discretized into 256 uniform bins across each DoFs’ action bounds $[y_{\min}^{i},y_{\max}^{i}]$ [34]. The $gripper$ state is binary, indicating its open or closed state. This control design presents a unique challenge for adversarial attacks, as finely divided bins result in minimal action discrepancies between neighboring bins (e.g., $\pm$0.007/bin). This means that even if an attack shifts the classification to an adjacent bin, the resulting action discrepancy has minimal impact on real-world performance.
@@ 3.2 Untargeted Action Discrepancy Attack
To exacerbate action discrepancies, we introduce the Untargeted Action Discrepancy Attack (UADA), which aims to maximize deviations in robot actions. This attack is based on the observation that larger robot actions usually correlate with intense physical movements, which, in turn, may amplify the potential for real-world hazards [28, 30, 29]. For UADA, the attack target is one or a combination of 7 DoFs, defined as $y^{I}_{gt}=\{y^{i}_{gt}|i\in[1,\dots,7]\}$. To define UADA’s objective, we first identify the most distant action $y_{adv}^{i}$, which maximizes the discrepancy from the $i$-th DoF ground truth action $y^{i}$. This action, for each DoF, corresponds to its own action bound $[y_{\min}^{i},y_{\max}^{i}]$ as:
$y_{adv}^{i}=\begin{cases}y_{\max}^{i},&\text{if }|y_{\max}^{i}-y^{i}_{gt}|\geq|y_{\min}^{i}-y^{i}_{gt}|\\ y_{\min}^{i},&\text{otherwise}\end{cases}.$ (2)
Instead of directly using $y_{adv}^{i}$ as the misclassification target, we introduce a soft attack objective to capture the discrepancy between actions, ensuring smooth gradient optimization and stable attack performance. Specifically, due to the physical action magnitude information contains in bin labels, we reweight the output probability $F(x)^{i}_{bins}\in\mathcal{R}^{{}_{1\times{bins}}}$ using normalized bin labels $y^{i}_{bins}=[\frac{1}{bins},\dots,1]$, where $\lfloor y^{i}_{bins}\rfloor$ and $\lceil y^{i}_{bins}\rceil$ correspond to the normalized values of $y_{\min}^{i}$ and $y_{\max}^{i}$, respectively. The reweighting process is defined as:
$y_{soft}^{i}=\sum_{bins=1}^{n}F(x+\delta)^{i}_{bins}\otimes y^{i}_{bins},$ (3)
where $\otimes$ denotes Hadamard Product, $\delta$ is the adversarial patch and $y_{soft}^{i}$ represents the $i$-th DoF’s soft action. Finally, the objective of UADA is to minimize the discrepancy between the soft and the most distant actions, which is:
$\mathcal{L}_{\text{UADA}}=\mathbb{E}_{(x,y)\sim\mathcal{X}}\sum_{i}^{I}(y_{soft}^{i}-y_{adv}^{i})^{2},$ (4)
UADA defines the objective function $\mathcal{L}_{\text{UADA}}$ tailored for attacking the robot’s action space, allowing adversarial patches to create significant, lasting disruptions in task performance.
@@ 3.3 Untargeted Position-aware Attack
We explore untargeted adversarial attacks in parallel by considering the positional dynamics within VLA models. In typical task execution, precise and directed movements of the end-effector towards designated goals are essential, with actions mapped in three-dimensional space. The positional DoFs, $A^{p}=[\Delta{P_{x}},\Delta{P_{y}},\Delta{P_{z}}]$ encapsulates the directional movements within 3D for effective robotic control.
Recognizing the importance of $A^{p}=DT(y^{p})$ in controlling the end-effector’s path, we introduce a position-aware attack to disrupt the intended movement trajectory. This type of adversarial vulnerability remains largely unexplored but has significant implications for task failure: the attack objective can steer the end-effector away from its intended path by introducing directional perturbations, amplifying errors and causing substantial trajectory distortions. ${}_{\!}$we define the Untargeted Position-aware Attack (UPA) objective as:
$\small\mathcal{L}_{\text{UPA}}=\mathbb{E}_{(x,y)\sim\mathcal{X}}\big{[}\alpha\frac{y^{p}_{adv}\cdot y^{p}}{\|y^{p}_{adv}\|\|y^{p}\|}+\beta\frac{1}{||y^{p}_{adv}-y^{p}||_{2}}\big{]},$ (5)
@@
3.4 Targeted Manipulation Attack
In addition to the aforementioned untargeted attacks, we also explore the adversarial vulnerabilities of VLA models with a targeted strategy. This attack aims to directly mislead the models into predicting specific trajectories, resulting in precise alterations. We design this attack since the targeted alterations can drive malicious behaviors or induce task failure. The objective of the targeted manipulation attack is:
$\small\mathcal{L_{\text{TMA}}}=\mathbb{E}_{(x,y)\sim\mathcal{X}}[CE(\mathcal{F}(x+\delta)^{I},y^{I}_{T})],$ (6)
where $y^{I}_{T}=\{y^{i}_{T}=t|i\in[1,\dots,7],t\in[y^{i}_{\min},y^{i}_{\max}]\}$ represents the adversarial target(s); and $CE(\cdot,\cdot)$ denotes the Cross-Entropy loss, which measures the error between the predicted state under adversarial perturbation and the desired target state. By steering the robot toward an adversarial target across successive time steps, our approach manipulates the trajectory and undermines task performance, ultimately leading to a substantial disruption of intended operations. Furthermore, manipulating any individual DoF or a combination thereof can also severely compromise task success, given that the errors in one dimension tend to propagate and magnify over time during execution (see Fig. 3).
3.5 A Novel Evaluation Metric:
Normalized Action Discrepancy
To assess the potential physical impact of adversarial patches, we introduce Normalized Action Discrepancy (NAD) as a fine-grained metric for evaluating discrepancies at the action level. NAD serves as a complementary measurement alongside the coarse-grained Failure Rate (FR) at the task level, providing a detailed analysis of deviations during action execution. To define the NAD, we first determine the maximum action discrepancy $d^{i}_{\text{max}}$, which represents the largest deviation that can occur within the allowed range for $i$-th DoF. This is computed by measuring the L1 distance between the ground truth $y^{i}_{gt}$ and its action bound $[y^{i}_{\min},y^{i}_{\max}]$ as:
$\small d_{\text{max}}^{i}(y)=\max\big{[}DT(|y^{i}_{gt}-y^{i}_{\min}|),DT(|y^{i}_{gt}-y^{i}_{\max}|)\big{]}.$ (7)
We next calculate the applied action discrepancy $d_{\text{applied}}^{i}(x,y)$ $=|DT(F(x)^{i})-DT(y^{i}_{gt})|$ to measure the deviation between the model’s output and ground truth. We define the NAD as the drifting ratio between $d_{\text{applied}}^{i}$ and $d_{\text{max}}^{i}$ as:
$\rm{NAD}=\frac{1}{I}\sum_{i}^{I}\frac{d^{i}_{\text{applied}}}{d^{i}_{\text{max}}},$ (8)
where $I\in[1,\dots,7]$ represents a single or a combination of DoF(s). In this way, NAD provides normalized measurement of action deviations, allowing for a consistent evaluation of adversarial impact across varying DoF(s).
4 Experiments
In this section, we first detail the implementation of our adversarial framework and baseline methods in §4.1. Next, we outline the experimental setup in §4.2 and present quantitative and qualitative results in both simulation and physical-world scenarios (§4.3). We then conduct diagnostic experiments (§4.4) to analyze the impact of key components. Additionally, we assess the robustness of our method against various defense mechanisms in §4.5 and conclude with a comprehensive systemic discussion in §4.6.
1:  Input: $\mathcal{X}$: dataset; $\delta$: patch; $\mathcal{L}_{o}$: attack objective; $\mathcal{F}$: VLA model; $\mathcal{T}(\cdot)$: transformations; $\phi,\psi$: transforma -tion params; T: attack steps; $k$: inner-loop steps.